# NeurIT: Pushing the Limit of Neural Inertial Tracking for Indoor Robotic IoT

### v1

@@ III-C Model Architecture
We propose TF-BRT to improve the learning outcome for IMU-based robot tracking. TF-BRT comprises six key components, as shown in Fig. 3: time-frequency input projection, block-recurrent attention module, convolution module, multi-head attention module, feedforward module, and output projection layer. The time-frequency input projection extracts relevant information from both the time and frequency domains, capturing hidden features related to the robot’s movement. The block-recurrent attention module incorporates RNN into the Transformer, allowing for pre-sequence extraction and contextual understanding of the current motion within the receptive field, which enables more accurate trajectory generation. To make the network convergence faster, the remaining four modules are interconnected in a residual way [23] with a pre-norm residual unit [24]. For the convolution module, a depthwise separable convolution [25] layer is sandwiched between the layernorm and dropout layer. Depthwise separable convolution can efficiently captures sequence’s local features. The multi-head attention module employs a multi-head self-attention layer positioned between the layernorm and dropout layers. Finally, the feedforward module utilizes two fully-connected layers to extract the final features. After all the processes are completed, the output projection layer effectively distills the robot’s velocity information from the high-dimensional space, resulting in the velocity sequence $\textbf{V}_{n}$. This design approach excels in handling long-time sequences and shows promising generalization abilities in various environments.
III-C1 Time-Frequency Input Projection
@@
Recent studies on time series prediction and classification have emphasized the importance of incorporating frequency domain information. It has been suggested that the distinct patterns of various time series can be expressed through the frequency domain for an enhanced performance [26, 27, 28]. One such approach, UniTS [26], incorporates Short-Time Fourier Transform (STFT) in its convolution kernel coefficient initialization and demonstrates improved performance in IMU motion recognition when compared to models relying solely on time domain information.
@@
Inspired by prior research, TF-BRT utilizes dual-channel inputs consisting of both time and frequency domain features. However, unlike UniTS and RF-Net [27], which use STFT and FFT as separate methods to extract frequency features, TF-BRT utilizes Discrete Cosine Transform (DCT) to extract frequency domain information. This choice avoids the need for complex number operations within neural networks.
@@
The two input channels are connected by projection layers, which extract hidden features in the two channels through the fully connected layer and map the IMU feature domain ($\mathbb{R}^{L\times D}$) to a larger domain ($\mathbb{R}^{L\times D_{h}}$), enabling subsequent networks to effectively learn and operate within the expanded high-dimensional feature space. $D_{h}$ represents the feature dimension in the following modules. The data after the projection layers are merged together to combine the extracted high-dimension features of the time domain and frequency domain. This operation can be written as $f_{\theta_{t}}(\textbf{X}_{n})+f_{\theta_{f}}({\rm{DCT}}(\textbf{X}_{n}))=\textbf{M}_{n}$, where $\theta_{t}$ and $\theta_{f}$ denote the projection layer parameters of time channel and frequency channel, respectively, and $\textbf{M}_{n}$ represents the merged hidden features.
@@
III-C2 Block-Recurrent Attention Module
Our model borrows the key framework of the block-recurrent transformer [22] to leverage the advantages of both RNN and self-attention mechanisms. Furthermore, we incorporate relative position encoding [29] as an additional component to augment the input features of the block-recurrent transformer. This encoding improves the representation of temporal order in the model. The block-recurrent attention module follows a pre-norm residual unit in a sandwich format. Specifically, the merged hidden features $\textbf{M}_{n}$ from the merge layer are first normalized using a layernorm operation. The normalized hidden features, or the input embedding for the next layer, can be expressed as follows:
$\textbf{E}_{n}={\rm{LayerNorm}}(\textbf{M}_{n})$ (3)
@@ III-C2 Block-Recurrent Attention Module
where $\rm{RPE}$ denotes the relative position encoding [29], which maps an order of indexes to the input sequences. After passing through a projection layer, $\textbf{H}_{n}$ is merged with $\textbf{E}_{n}$, and the output embedding is finally obtained through a multilayer perceptron (MLP) layer based on residual format:
@@
Fig. 4b gives the workflow for the next state $\textbf{H}_{n+1}$. The only difference is that it utilizes LSTM gate. In addition, RNN structure allows the state tensor to selectively merge the current and previous information, providing essential information for future prediction. It should be noted that all of these tensors share the space of $\mathbb{R}^{L\times D_{h}}$. Finally, $\textbf{E}_{n+1}$ is sent to the next module after the dropout layer.
III-D Multi-Loss Learning
To improve TF-BRT’s regression performance for estimating moving velocities and heading orientations, we employ a multi-loss learning strategy. The multi-loss function consists of three components: velocity loss $\mathcal{L}_{v}$, position loss or integrated velocity loss $\mathcal{L}_{p}$, and orientation loss $\mathcal{L}_{o}$.
Velocity Loss ($\mathcal{L}_{v}$): With a predicted velocity sequence $\textbf{V}_{n}=\{v_{1},v_{2},...,v_{L}\}$, a mean square loss is utilized to measure the difference between the ground truth velocity $\textbf{V}_{n}^{g}=\{v_{1}^{g},v_{2}^{g},...,v_{L}^{g}\}$. The velocity loss function can be written as $\mathcal{L}_{v}=\sqrt{\frac{1}{L}\sum_{i=1}^{L}(v_{i}-v_{i}^{g})^{2}}$.
Position Loss ($\mathcal{L}_{p}$): The predicted velocity is integrated to obtain the predicted moving trajectory $\textbf{P}_{n}$. By using $\mathcal{L}_{p}$, we aim to further reduce the cumulative error based on the learned velocity characteristics. The equation for $\mathcal{L}_{p}$ is $\mathcal{L}_{p}=\sqrt{\frac{1}{L}\sum_{i=1}^{L}(p_{i}-p_{i}^{g})^{2}}$.
Orientation Loss ($\mathcal{L}_{o}$): The major prediction errors come from direction deviations, especially at low speeds. For instance, even if the velocity loss is comparable to that at high speeds, the actual orientation error can be substantial due to the division of velocity error by the velocity modulus. To address this issue, we incorporate orientation loss to improve the accuracy of the predicted velocity in the travel direction. The definition of $\mathcal{L}_{o}$ is as follows: $\mathcal{L}o=\sqrt{\frac{1}{L}\sum{i=1}^{L}(\frac{v_{i}}{||v_{i}||_{2}}-\frac{v_{i}^{g}}{||v_{i}^{g}||_{2}})^{2}}$.
@@ III-D Multi-Loss Learning
$\mathcal{L}_{total}=\frac{\sigma_{\mathcal{L}_{v}}}{\mu_{\mathcal{L}_{v}}}\mathcal{L}_{v}+\frac{\sigma_{\mathcal{L}_{p}}}{\mu_{\mathcal{L}_{p}}}\mathcal{L}_{p}+\frac{\sigma_{\mathcal{L}_{o}}}{\mu_{\mathcal{L}_{o}}}\mathcal{L}_{o}$ (6)
where $\mu$ and $\sigma$ denote the mean and standard deviation of the respective loss function. The reason for applying both velocity loss and position loss is that $\mathcal{L}_{v}$ corrects the velocity vector on each frame, while $\mathcal{L}_{p}$ corrects it on a time window scale. By combining local and global information, the model can better learn the characteristics of the motion. A more detailed ablation analysis can be found in §VI-C later.
III-E Data Augmentation
@@
where ${\Delta\textbf{q}}$ represents the random rotation in quaternion. By taking this data augmentation method, the moving orientation will distribute more evenly on the floor plane, thus making the model more robust. We conduct an ablation study of the data augmentation method in §VI-D, which further verifies our consumption.
IV Implementation
Data Collection: We build a customized robot platform for data collection, which captures raw IMU data as NeurIT input, visual-inertial tracking results from a commercialized device (Tango Phone [^1] https://www.lenovo.com/il/en/tango/) for comparison, and the ground truth data from a LiDAR-inertial navigation system based on FAST-LIO [2]. We choose the Tango phone and FAST-LIO for specific reasons. The Tango phone has been used as a ground truth in inertial tracking datasets like RIDI and RONIN. FAST-LIO is known for its minimal drift in indoor environments.
Fig. 5a illustrates the system setup. The main component of the system is an Intel NUC with an i7-1260P CPU. By obtaining point clouds through LIVOX MID360 (FOV: 360°*59°) [^2] FOV: 360°*59°; Point Cloud Density: 40-line. https://www.livoxtech.com/mid-360 and the built-in IMU, the FAST-LIO algorithm, which runs on the NUC, can accurately determine the indoor location of the robot. HiPNUC CH110 9-axis IMU is set on the left side, which records the raw data for model input. The Tango Phone set in the front is used to collect visual-inertial tracking results for comparison in §VI-F. In addition, we add AVIA (FOV: 70.4°*77.2°) [^3] FOV: 70.4°*77.2°; https://www.livoxtech.com/avia , a narrow-FOV LiDAR, to demonstrate the fact that LiDAR-inertial tracking systems cannot work well in simple and plain indoor environments. The robot’s territory is AgileX Tracer-Mini, whose size is $40cm\times 40cm\times 26cm$ with a maximum speed of 1.5m/s.
@@ IV Implementation
Figure 6: Topview of the locations of three buildings.
Model Training: Models are implemented using PyTorch 1.7.1 and Cuda 11 in Python. The experiments are run using NVIDIA 4090 with 24GB GPU memory. We use the same training settings for all baselines, with a batch size of 72, an initial learning rate of 0.0003, and the ADAM optimizer while reducing the learning rate by a factor of 0.75 if the validation loss does not decrease in 10 epochs.
V Experiments
In the part of experimental analysis, the model proposed in this paper is analyzed comprehensively under different data sets and different indicators. Furthermore, it includes a comparison of the model with the visual positioning algorithm in complex indoor scenes, as well as an ablation study of the model itself. The study also involves the system setup and the evaluation of the real-time system.
@@ V Experiments
V-A Datasets
Due to the lack of IMU datasets on robots subject, we build our own dataset to verify the effectiveness of TF-BRT. We split the dataset into four parts, i.e. training, validation, test-seen and test-unseen set. More details are described below:
NeurIT: We build our own NeurIT dataset using the customized robotic platform in four floors across three buildings, which feature different layouts, environments, and floor conditions, as shown in Fig. 5b. The locations of these three buildings are illustrated in Fig. 6. We collect the training, validation, and test-seen sets in two floors of Building A, and build the test-unseen set in Building B and C. During data collection, the robot moves at varying speeds up to the maximum value (1.5m/s). The dataset contains 110 sequences, totaling around 15 hours of tracking data that corresponds to a travel distance of about 33.7 km. Each sequence of data lasts 6$\sim$10 minutes. The ratio of the training set, validation set, test-seen set, and test-unseen set is 15:3:3:4.
RIDI: RIDI contains walking data of 10 human participants. We follow the default split setting in this dataset, which is 5:1:2:1.
RONIN: RONIN involves 100 human subjects. It is divided according to the default setting, i.e., 69:16:35:32. It should be noted only 50% of the dataset has been released.
During the training stage, we divide the NeurIT and RIDI according to the segmentation size of $200\times 15$ and the step size of $400$, and slides $15$ times according to the window size of $200$ on each segment sequence. We partition RONIN dataset according to the partition size of $400\times 20$ and the step size of $1000$, and slide $20$ times on each partition sequence according to the window size of $400$. During the test stage, we use the same window size and step size of $200$ for NeurIT and RIDI, while a window size and step size of 400 for RONIN.
V-B Baselines
To comprehensively evaluate NeurIT, we compare it with six state-of-the-art baseline approaches.
@@ V-B Baselines
  • Extended Kalman Filter (EKF): EKF [7] filters out Gaussian-like noise in the IMU data and combines it with the magnetometer to correct the orientation error.
@@
It should be noted that Transformer-based solutions, such as CTIN [14] and A2DIO [15], even though they achieve promising results in inertial tracking, the model cannot be reproduced without implementation details and the source code. In this case, we do not include them as baseline models. We will evaluate the performance comparison with Transformer-based methodology in the ablation study §VI-C.
V-C Evaluation Metrics
  • Absolute Trajectory Error (ATE): ATE [34] calculates the root mean square error (RMSE) of the predicted trajectory and the ground truth trajectory, which is given as $\sqrt{\frac{1}{n}\sum_{i}\|p_{i}-p^{g}_{i}\|_{2}^{2}}$.
  • Relative Trajectory Error (RTE): RTE [34] measures the RMSE between the predicted trajectory and the ground truth over a time-interval. RTE can be calculated as $\sqrt{\frac{1}{n}\sum_{i}\left\|(p_{i+\Delta t}-p_{i})-(p^{g}_{i+\Delta t}-p^{g}_{i})\right\|_{2}^{2}}$, where $\Delta t$ is set to 60s in our evaluation.
  • Position Drift Error (PDE): PDE [14] computes the final position drift over the whole sequence. It can be expressed as $\|p_{N}-p^{g}_{N}\|_{2}/{D}$, where $N$ is the last index of the sequence, and $D$ denotes the ground truth trajectory length.
  • Absolute Yaw Error (AYE): AYE [35] measures the RMSE between the heading orientations between ground truth and predicted trajectories, i.e., $\sqrt{\frac{1}{n}\sum_{i}\left\|\frac{v_{i}}{\|v_{i}\|_{2}}-\frac{v^{g}_{i}}{\|v^{g}_{i}\|_{2}}\right\|_{2}^{2}}$.
@@ V-C Evaluation Metrics
Metric NDI EKF RONIN IONet UniTS RF-Net
@@
ATE 7.93 7.65 3.59 3.26 2.71 5.35 1.53 43.56% ↑ 1.01 62.89% ↑
RTE 10.01 9.61 2.19 1.58 1.37 2.09 1.01 26.18% ↑ 0.58 57.31% ↑
PDE 0.029 0.040 0.026 0.027 0.021 0.048 0.010 52.88% ↑ 0.007 67.26% ↑
Seen AYE 123.09 124.77 63.77 46.23 121.41 99.76 37.90 18.02% ↑ 35.99 22.15% ↑
ATE 9.24 8.96 4.26 3.45 2.90 4.12 1.77 39.11% ↑ 1.14 60.89% ↑
RTE 11.24 11.10 2.42 1.87 1.74 2.24 1.61 7.09% ↑ 0.76 56.42% ↑
PDE 0.020 0.036 0.035 0.027 0.023 0.035 0.013 36.81% ↑ 0.009 55.70% ↑
NeurIT Dataset Unseen AYE 132.90 134.11 51.66 40.10 119.55 89.36 33.58 16.25% ↑ 32.14 19.86% ↑
Figure 7: CDF of ATE and RTE on NeurIT Test Dataset
Figure 8: Examples of Predicted Trajectories of NeurIT
VI Evaluation
VI-A Overall Performance
Evaluation on NeurIT Dataset: Tab. II shows the overall evaluation on NeurIT dataset. As seen, NeurIT produces remarkable accuracy on both seen and unseen test data, with a PDE of nearly 0.08, which implies an average error of only 0.8cm for every 1m traveled. NeurIT significantly reduces the average ATE by 62% over the best baseline approach. The extremely low values of RTE indicate a tiny drift of around 0.6% (0.6m drift over 1 minute of tracking). While UniTS (best baseline) can only achieve RTE results of approximately 1.5m, which is more than double. Furthermore, thanks to the multi-loss learning, NeurIT observes the improvement in AYE. NeurIT reduces the heading error by around 20%, from 43° (IONet) to 33°, indicating effective orientation error correction. Even without integrating magnetometers as input features, NeurIT outperforms the best baseline approach by an average of 30% across four metrics. For instance, by incorporating differential magnetometers, NeurIT achieves much greater performance enhancements, especially the results of RTE, which are almost reduced by 50%. These outcomes underscore the effectiveness of our sensor-fusion design.
To provide a more comprehensive visualization of the results, we include the cumulative distribution function of ATE and RTE in Fig. 7. The plots demonstrate that NeurIT surpasses all baseline models in terms of accuracy and reliability, particularly in unseen test sets. Even in the worst case, NeurIT outperforms the best-case performance of the other baseline models. We visualize some examples of the predicted trajectories in Fig. 8, to illustrate the stable tracking ability of NeurIT. The visualizations depict that the tracking trajectories generated by NeurIT closely match the ground truth trajectories. In contrast, other methodologies exhibit significant deviations, indicating their inability to track the robot’s movement.
An additional comparison among the baseline models also demonstrates the shortcomings of using RNN or Transformers only, for inertial tracking. For instance, UniTS almost performs the best among all baseline models, however, it suffers from significant AYE and also fails in certain cases, resulting in long-tail drift errors. In contrast, by combining the advantages of RNN and Transformer, NeurIT can achieve state-of-the-art accuracy and robustness.
@@ VI-A Overall Performance
Metric NDI EKF RONIN IONet UniTS RF-Net

### v2

@@ IV-C Model Architecture
We construct TF-BRT based on advanced neural network architectures to improve the learning outcome for IMU-based robot tracking. TF-BRT comprises six key components, as shown in Fig. 3: time-frequency input projection, block-recurrent attention [36] module, convolution [37] module, multi-head attention [38] module, feedforward module, and output projection layer. The time-frequency input projection captures relevant features from both domains, revealing latent patterns in the robot’s motion. The block-recurrent attention module integrates RNNs into the Transformer, enabling pre-sequence modeling and contextual understanding within the receptive field for more accurate trajectory prediction. To make the network convergence faster, the remaining four modules are interconnected in a residual way [39] with a pre-norm residual unit [40]. For the convolution module, a depthwise separable convolution [41] layer is sandwiched between the layernorm and dropout layer. Depthwise separable convolution can efficiently capture a sequence’s local features. The multi-head attention module employs a multi-head self-attention layer positioned between the layernorm and dropout layers. Finally, the feedforward module utilizes two fully-connected layers to extract the final features. After all the processes are completed, the output projection layer effectively distills the robot’s velocity information from the high-dimensional space, resulting in the velocity sequence $\textbf{V}_{n}$. This design approach excels in handling long-time sequences and shows promising generalization abilities in various environments.
IV-C1 Time-Frequency Input Projection
@@
Recent studies on time series prediction and classification have emphasized the importance of incorporating frequency domain information. It has been suggested that the distinct patterns of various time series can be expressed through the frequency domain for an enhanced performance [42, 43, 44]. One such approach, UniTS [42], improves IMU motion recognition by using Short-Time Fourier Transform (STFT) to initialize the convolution kernels, surpassing models that rely only on time-domain data.
@@
Inspired by prior research, TF-BRT utilizes dual-channel inputs consisting of both time and frequency domain features. However, unlike UniTS and RF-Net [43], which use STFT and FFT as separate methods to extract frequency features, TF-BRT utilizes the Discrete Cosine Transform (DCT) to extract frequency domain information, avoiding the need for complex number operations in the network.
@@
The two input channels are connected by projection layers, which extract hidden features in the two channels through the fully connected layer and map the IMU feature domain ($\mathbb{R}^{L\times D}$) to a larger domain ($\mathbb{R}^{L\times D_{h}}$), enabling subsequent networks to effectively learn and operate within the high-dimensional feature space. $D_{h}$ represents the feature dimension in the following modules. The data after the projection layers are merged to combine the extracted high-dimensional features of the time domain and frequency domain. This operation can be written as $f_{\theta_{t}}(\textbf{X}_{n})+f_{\theta_{f}}({\rm{DCT}}(\textbf{X}_{n}))=\textbf{M}_{n}$, where $\theta_{t}$ and $\theta_{f}$ denote the projection layer parameters of time channel and frequency channel, respectively, and $\textbf{M}_{n}$ represents the merged hidden features.
@@
IV-C2 Block-Recurrent Attention Module
Our model borrows the key framework of the block-recurrent transformer [36] to leverage the advantages of both RNN and self-attention mechanisms. Furthermore, we incorporate relative position encoding [45] as an additional component to augment the input features of the block-recurrent transformer. This encoding improves the representation of temporal order in the model. The block-recurrent attention module follows a pre-norm residual unit in a sandwich format. Specifically, the merged hidden features $\textbf{M}_{n}$ from the merge layer are first normalized using a layernorm operation. The normalized hidden features, or the input embedding for the next layer, can be expressed as follows: $\textbf{E}_{n}={\rm{LayerNorm}}(\textbf{M}_{n})$.
@@ IV-C2 Block-Recurrent Attention Module
where $\rm{RPE}$ denotes the relative position encoding [45]. After passing through a projection layer, $\textbf{H}_{n}$ is merged with $\textbf{E}_{n}$, and the output embedding is finally obtained through a multilayer perceptron (MLP) layer in a residue manner:
@@
Fig. 4b gives the workflow for the next state $\textbf{H}_{n+1}$. The only difference is that it utilizes LSTM gate. In addition, RNN structure allows the state tensor to selectively merge the current and previous information, providing essential information for future prediction. Finally, $\textbf{E}_{n+1}$ is sent to the next module after the dropout layer.
IV-D Multi-Loss Learning
To improve TF-BRT’s regression performance for estimating moving velocities and heading orientations, we employ a multi-loss learning strategy. The multi-loss function consists of three components: velocity loss $\mathcal{L}_{\mathbf{v}}$, position loss $\mathcal{L}_{\mathbf{p}}$, and orientation loss $\mathcal{L}_{\mathbf{o}}$.
Velocity Loss ($\mathcal{L}_{\mathbf{v}}$): With a predicted velocity sequence $\textbf{V}_{n}=[\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{L}]$, a mean square loss is utilized to measure the difference between the ground truth velocity $\textbf{V}_{n}^{g}=[\mathbf{v}_{1}^{g},\mathbf{v}_{2}^{g},...,\mathbf{v}_{L}^{g}]$. The velocity loss function can be written as $\mathcal{L}_{\mathbf{v}}=\sqrt{\frac{1}{L}\sum_{i=1}^{L}(\mathbf{v}_{i}-\mathbf{v}_{i}^{g})^{2}}$.
Position Loss ($\mathcal{L}_{\mathbf{p}}$): The predicted velocity is integrated to obtain the predicted moving trajectory $\textbf{P}_{n}$. By using $\mathcal{L}_{\mathbf{p}}$, we aim to further reduce the cumulative error based on the learned velocity characteristics. The equation for $\mathcal{L}_{\mathbf{p}}$ is $\mathcal{L}_{\mathbf{p}}=\sqrt{\frac{1}{L}\sum_{i=1}^{L}(\mathbf{p}_{i}-\mathbf{p}_{i}^{g})^{2}}$.
Orientation Loss ($\mathcal{L}_{\mathbf{o}}$): The major prediction errors come from direction deviations, especially at low speeds. For instance, even if the velocity loss is comparable to that at high speeds, the actual orientation error can be substantial due to the division of velocity error by the velocity modulus. To address this issue, we incorporate orientation loss to improve the accuracy of the predicted velocity in the travel direction. The definition of $\mathcal{L}_{\mathbf{o}}=\sqrt{\frac{1}{L}\sum_{i=1}^{L}(\frac{\mathbf{v}_{i}}{||\mathbf{v}_{i}||_{2}}-\frac{\mathbf{v}_{i}^{g}}{||\mathbf{v}_{i}^{g}||_{2}})^{2}}$.
@@ IV-D Multi-Loss Learning
$\mathcal{L}_{total}=\frac{\sigma_{\mathcal{L}_{\mathbf{v}}}}{\mu_{\mathcal{L}_{\mathbf{v}}}}\mathcal{L}_{\mathbf{v}}+\frac{\sigma_{\mathcal{L}_{\mathbf{p}}}}{\mu_{\mathcal{L}_{\mathbf{p}}}}\mathcal{L}_{\mathbf{p}}+\frac{\sigma_{\mathcal{L}_{\mathbf{o}}}}{\mu_{\mathcal{L}_{\mathbf{o}}}}\mathcal{L}_{\mathbf{o}}$ (4)
where $\mu$ and $\sigma$ denote the mean and standard deviation of the respective loss function. The reason for applying both velocity loss and position loss is that $\mathcal{L}_{\mathbf{v}}$ corrects the velocity vector on each frame, while $\mathcal{L}_{\mathbf{p}}$ corrects it on a time window scale. By combining local and global information, the model can learn the characteristics of the motion better. A detailed ablation analysis can be found in §VII-C.
IV-E Data Augmentation
@@
where ${\Delta\textbf{q}}$ represents the random rotation in quaternion. This data augmentation helps distribute movement orientations more evenly on the floor plane, improving model robustness. An ablation study in §VII-D further supports its effectiveness.
@@
V Implementation
We build a customized robotic platform for data collection, which captures raw IMU data as NeurIT input, visual-inertial tracking results from a commercialized device (Tango Phone  [^1] https://www.lenovo.com/il/en/tango/ ) for comparison, and the ground truth data from a LiDAR-inertial tracking system implemented with FAST-LIO [2]. We choose the Tango phone and FAST-LIO for specific reasons. Tango phone has been used as the ground truth in inertial tracking datasets like RIDI and RONIN. FAST-LIO is known for its minimal drift in indoor environments.
Fig. 5a illustrates the system setup. The core of the system is an Intel NUC with an i7-1260P CPU. It runs the FAST-LIO algorithm using point clouds from the LIVOX MID360222FOV: 360°×59°; Point Cloud Density: 40-line. https://www.livoxtech.com/mid-360 and its built-in IMU to determine the robot’s indoor position. A HiPNUC CH110 9-axis IMU is mounted on the left side to collect raw data for model input. A Tango phone, placed at the front, provides visual-inertial tracking results for comparison in §VII-G. In addition, we add AVIA  [^3] FOV: 70.4°*77.2°; https://www.livoxtech.com/avia , a narrow-FOV LiDAR, to demonstrate the fact that LiDAR-inertial tracking systems cannot work well in plain indoor environments. The robot’s territory is AgileX Tracer-Mini, whose size is $40\rm{cm}\times 40\rm{cm}\times 26\rm{cm}$ with a maximum speed of 1.5m/s.
@@ V Implementation
Figure 6: Topview of the locations of three buildings.
VI Experiments
We evaluate NeurIT across diverse datasets and metrics. To test its robustness, we conduct experiments in large urban spaces, such as shopping malls and concert halls. We also compare it with a visual-inertial method indoors, perform ablation studies to analyze component contributions, and assess real-time performance to validate practical deployment.
VI-A Datasets
Due to the lack of IMU datasets on robot subjects, we built our dataset to verify the effectiveness of NeurIT. We split the dataset into four parts, i.e., training, validation, test-seen, and test-unseen sets. More details are described below:
NeurIT Dataset: We built the NeurIT dataset using the customized robotic platform on four floors across three buildings, which feature different layouts, environments, and floor conditions, as shown in Fig. 5b. The locations of these three buildings are illustrated in Fig. 6. We collect the training, validation, and test-seen sets in Building A, and build the test-unseen set in Building B and C. During data collection, the robot moves at varying speeds up to the maximum value (1.5m/s).
The dataset contains 110 sequences, totaling around 15 hours of tracking data that corresponds to a travel distance of around 33.7 km. Each sequence of data lasts 6$\sim$10 minutes. The ratio of the training set, validation set, test-seen set, and test-unseen set is 15:3:3:4.
RIDI: RIDI contains the walking data of 10 human participants. We follow the default split setting in this dataset, which is 5:1:2:1.
RONIN: RONIN involves 100 human subjects. It is divided according to the default setting, i.e., 69:16:35:32. It should be noted that only 50% of the dataset has been publicly released.
During the training stage, we divide the NeurIT and RIDI according to the segmentation size of 200×15200$\times$ 15200 × 15 and the step size of 400400400, and slide 151515 times according to the window size of 200200200 on each segment sequence. We partition the RONIN dataset into segments of size 400×20400$\times$ 20400 × 20 with a step size of 1000. Each segment is then processed using a sliding window of size 400, applied 20 times per sequence. During testing, we use a window size and step size of 200 for both NeurIT and RIDI, and 400 for RONIN..
VI-B Baselines
To comprehensively evaluate NeurIT, we compare it with seven state-of-the-art baseline approaches.
@@ VI-B Baselines
  • Extended Kalman Filter (EKF): EKF [7] filters out Gaussian-like noise in the IMU data and combines it with the magnetometer to correct the orientation error. We use the python implementation [49] in this work.
@@
  • LLIO: LLIO [24] utilizes multiple MLP layers in a residual way to predict the trajectories for pedestrians. The input features consist of accelerations and gyroscopes.
@@
It should be noted that Transformer-based solutions, such as CTIN [14] and A2DIO [15], even though they achieve promising results in inertial tracking, the model cannot be reproduced without implementation details and the source code. In this case, we do not include them as baseline models. We will evaluate the performance comparison with Transformer-based methodology in the ablation study §VII-C.
@@
Metric NDI EKF RONIN IONet LLIO UniTS RF-Net
@@
ATE 7.926 7.651 3.590 3.262 3.061 2.712 5.353 1.539 43.26% ↑ 1.010 62.76% ↑
RTE 10.005 9.608 2.190 1.584 1.592 1.370 2.085 1.083 20.95% ↑ 0.580 57.67% ↑
PDE 0.029 0.040 0.026 0.027 0.024 0.021 0.048 0.010 51.31% ↑ 0.007 67.38% ↑
Seen AYE 123.086 124.774 63.770 46.228 58.618 121.414 99.765 27.544 40.42% ↑ 35.990 22.15% ↑
ATE 9.238 8.958 4.260 3.446 3.167 2.904 4.125 1.755 39.56% ↑ 1.140 60.75% ↑
RTE 11.241 11.099 2.420 1.874 1.894 1.736 2.236 1.538 11.38% ↑ 0.760 56.21% ↑
PDE 0.020 0.036 0.035 0.027 0.024 0.023 0.035 0.013 32.88% ↑ 0.009 55.01% ↑
NeurIT Unseen AYE 132.905 134.107 51.660 40.101 60.501 119.548 89.364 22.950 42.77% ↑ 32.140 19.86% ↑
@@
Metric NDI EKF RONIN IONet LLIO UniTS RF-Net